# EVINCE: Optimizing Adversarial LLM Dialogues via Conditional Statistics and Information Theory

### v1

@@ 3.3 Maxims with Theoretical Foundations
Progress towards the optimality goal is guided and measured by metrics introduced in Section 3.1. This section explains how these metrics can be used in complementary ways to facilitate proper trade-offs between diversity and convergence, exploration and exploitation, and several other factors. In the EVINCE algorithm presented in Figure 1, we have annotated the steps to which these four maxims are applied.
Maxim #1: Orchestrate Two Equally Competent LLMs in Structured Debate:
@@
How? Choosing LLMs with comparable performance on a shared validation set, a balanced debate can be ensured. Suitable models include GPT-4, Claude, and Gemini. Conditioning different instances of the same LLM to support opposing stances on a subject matter can also be effective due to the theoretical justification of in-context learning with conditional Bayesian statistics Xie et al. (2021).
Maxim #2: Encourage the Accurate Rather Than the “Popular” Prediction:
Typically, LLMs, with their maximum likelihood next-token prediction objective, tend to favor the most popular predictions. By conditioning LLMs within specific contexts, we can prioritize specific stance over popularity, mitigating confirmation biases.
How? Using the proxy metrics in Table 1, EVINCE dynamically adjusts the “contentiousness” level in debates. These metrics quantify agreement, diversity, and mutual information, promoting productive information exchange and enhancing prediction quality.
Maxim #3. Evaluating the Convergence Rate of the Predictions Across the Iterations:
This aspect focuses on measuring how quickly and effectively the predictions from the LLMs converge over successive rounds, assessing the efficiency of the debate and aggregation mechanisms.
How? Convergence is assessed by measuring mutual information and using proxy metrics such as Wasserstein distance and cross entropy. When the mutual information is high or the similarity between predictions is high, the debate is considered to be converging.
@@
Weighting the probability distributions from two LLMs based on diverse probabilistic insights and argument quality.
How? Following these three maxims:
  • Maxim #4.1 Prediction Reliability: Estimate the reliability of predictions using entropy-based measures to quantify uncertainty and information content. Typically, lower entropy indicates higher confidence in a prediction, suggesting higher reliability.
  • Maxim #4.2 Argument Quality: Evaluate the quality of supporting arguments using techniques inspired by the Socratic method. This includes identifying logical fallacies, assessing the relevance and credibility of evidence.
  • Maxim #4.3 Aggregation: Employ a weighted aggregation method, such as a Bayesian model to combine weighted predictions accounting for both probabilistic insights and the quality of supporting arguments.
@@ 3.4 Entropy Duality Theorem (EDT)
Please see Table 3. ∎
[Proof]:
[…]
Step 6: Conclude the proof. By combining the prediction distributions of LLMA and LLMB, with one having high information entropy and the other low, we achieve an optimal balance that maximizes the lower bound of the combined information entropy. This balance between exploration (high entropy) and exploitation (low entropy) optimizes overall prediction accuracy and robustness, as proved through information theory and the properties of entropy. Thus, the theorem is established. $\square$
Table 3: Proof of Entropy Dual Theorem (EDT)
4 Empirical Study
This empirical study investigates the application of EVINCE to disease diagnosis, leveraging large language models (LLMs) as diagnostic tools. We aim to validate the following three hypotheses:
@@ 4 Empirical Study
    Contentiousness & Prediction Quality: Initial LLM disagreement (measured by Wasserstein distance) increases with higher initial contentiousness but decreases as debate progresses. Individual LLM prediction uncertainty (Shannon entropy) will follow a similar pattern.
@@
    EDT Effectiveness & Confusion Matrices: LLM pairs following the Entropy Duality Theorem (EDT) will have complementary error patterns, leading to higher combined prediction accuracy than non-EDT pairs.
@@
    EVINCE & Historical Misdiagnoses: EVINCE, applied to real-world data, will improve diagnostic accuracy and identify potential misdiagnoses or ambiguities within the ground truth.
@@
Figure 2: Pre-/post-debate accuracy on all patients on all diseases shows EVINCE helps
Problem Statement:
Given a set of symptoms $S$ and a context $\kappa$, the goal is to predict a probability distribution of top-$k$ diseases over a set of $C$ possible diseases. This is represented as $P=\text{LLM}(S,\kappa)$, where each LLM generates top-$k$ predictions on $C$ ($k\leq|C|$) based on the input symptoms $S$ and context $\kappa$.
$P=\left(p(\text{top}\,1\,\text{to}\,k\in C\mid S,\kappa)\right).$
Context $\kappa$ is where dual entropy is adjusted through three parameters: temperature, the value of $k$ in top-$k$, and the contentious level $\Delta$. A distribution tends to have high entropy when all three parameters are set high, and vice versa.
Resources, Dataset & Data Preparation:
Our study utilizes a dataset from Kaggle Patil (2020), comprising 4,921 patient records. Each record includes the diagnosed disease and up to 17 symptoms (e.g., fever, cough, fatigue, itchiness, difficulty breathing). After removing duplicates, we obtain 304 unique diagnostic instances spanning 40 diseases. (The refined dataset is provided as supplementary data.) Each instance serves as a test case where EVINCE leverages the inherent knowledge of LLMs (GPT-4, Gemini, and Claude3) without training them through few-shot techniques on this specific dataset. Our computing resources are sponsored by Azure through a Stanford grant.
Evaluation:
We assess prediction quality using the top-k Mean Reciprocal Rank (MRR). If one of the top-k predicted diseases matches the ground truth diagnosis, the score is the reciprocal of its rank (1 for the top prediction, 1/2 for the second, 1/3 for the third, etc.). If none of the top-k predictions are correct, the score is 0.
4.1 Study #1: Post vs. Pre-Debate Accuracy
We employed GPT-4, Gemini, and Claude3 to perform independent disease predictions on 304 patient instances, then used EVINCE to pair them and evaluate performance gain.
Experimental Setup:
We set $k=5$ for both LLM agents, with one agent at high temperature and the other at low temperature. The contentiousness level was set very high ($\Delta=0.9$ out of 1) to encourage significant cross entropy. Setting $k=5$ ensures some minimal common ground, fostering meaningful interaction. High contentiousness promotes counterarguments and information exchange.
Pre- and Post-Debate Evaluation:
We conducted two sets of experiments:
1. Constrained Prediction (Baseline): We limited disease predictions to the 40 labels in the dataset, mimicking common supervised learning assumptions. This yielded high accuracy (95-97%) but is unrealistic for real-world diagnosis where general practitioners consider all possibilities. This constraint highlights LLMs’ flexibility, as they’re less prone to overfitting erroneous labels (further discussed in subsequent studies).
2. Unconstrained Prediction: We removed the label constraint to better simulate real-world conditions. All 304 patient cases yielded stable results across GPT-4, Gemini-3, and Claude-3, with a standard deviation of just 1.5%. Prior to debate (light blue bars in Figure 2), GPT-4 led in accuracy (82.8%), followed by Gemini (80.3%) and Claude (79.5%).
EVINCE Performance:
Implementing EVINCE with GPT-4 and Claude-3 pairing and GPT-4 and Gemini-3 pairing consistently improved accuracy by 4-5 percentage points (green bars in Figure 2). The GPT-4 and Claude-3 pairing achieved 87.5% accuracy (Figure 2(a)), rivaling state-of-the-art clinical performance like the REFUEL algorithm Peng et al. (2018).
Discussion:
The remaining 12.5% inaccuracy for the GPT-Claude pairing might not be solely attributable to EVINCE. Considering the potential 11% US misdiagnosis rate reported by Johns Hopkins Newman-Toker et al. (2023b), this discrepancy could indicate mislabeled data in the original dataset. This presents a groundbreaking opportunity: EVINCE could potentially identify and correct errors in existing datasets, a concept we explore further in Section 4.3.
4.2 Study #2: Confusion vs. Opportunities
@@ Discussion:
Two primary factors contribute to EVINCE’s improved diagnostic accuracy: (1) structured debates with reasoning encourage LLMs to explore alternative diagnoses in both breadth and depth, leading to more comprehensive analysis and decision-making (see Appendices B and C); and (2) pairing high- and low-entropy LLMs balances exploratory diversity with exploitative stability, resulting in more robust and high-quality decisions, as demonstrated in this second study.
@@
We use confusion matrices to analyze two LLMs’ performance in diagnosing Hepatitis types A to E (Fig. 3):
  • GPT-4 shows limited accuracy, particularly for types C and D, achieving only 50% accuracy for types A and B.
  • Claude exhibits a wider spread of predictions across all Hepatitis types.
These matrices highlight how Claude’s flexibility in exploring diverse diagnostic hypotheses aids the debate process. Claude’s initial uncertainty (high entropy) brings new information, potentially challenging and correcting GPT-4’s more confident (low entropy) predictions. This dynamic interplay exemplifies EVINCE’s balance between exploration and exploitation, leading to more accurate and comprehensive diagnoses.
Observations from Information Metrics:
  • Entropy Stabilization: Figure 4(a) shows entropy levels of both LLMs stabilizing after three debate rounds, indicating convergence towards a similar, stable entropy state.
  • Wasserstein Distance Improvement: Figure 4(b) demonstrates consistent improvement in Wasserstein distance (WD) between the two models’ predictions over successive rounds.
  • Mutual Information Increase: Figure 4(c) reveals a 14% improvement in normalized mutual information (MI) between GPT-4 and Claude’s prediction distributions, suggesting increased shared information throughout the debate.
  • Convergence of Divergence Metrics: Figure 5 shows consistent convergence of all divergence metrics.
@@
EVINCE demonstrates a 5% higher accuracy rate in diagnosing specific types of liver diseases compared to a baseline approach (Figure 2(a)), underscoring its capability to handle complex diagnostic scenarios effectively.
Figure 5: Convergence of all metrics
4.3 Study #3: Ground-Truth Remediation
This study illustrates how EVINCE can identify potential misdiagnoses, explain the reasoning behind them, and recommend corrective actions. Traditionally, machine learning scientists rely on labeled data as “ground truth.” However, as evidenced by research like that of Newman-Toker et al. (2021) Newman-Toker et al. (2023a) from Johns Hopkins, misdiagnosis is a widespread issue in healthcare systems globally. These erroneous diagnoses, often treated as ground truth, can be perpetuated by supervised learning algorithms, exacerbating the problem within the healthcare system.
@@ 4.3 Study #3: Ground-Truth Remediation
In the debate scenario detailed in Appendix C, where Jaundice is the ground truth diagnosis, Figure 6(a) illustrates initial differences between GPT-4 and Claude’s predictions. Jaundice is absent in GPT-4’s top-5 (with 0% in red), while ranked third by Claude. Although Claude influences GPT-4 to include Jaundice in its third prediction in the second round, subsequent rounds see both LLMs drop Jaundice to the fourth position of 10%.
Meanwhile, Hepatitis A, initially GPT-4’s top prediction (30% in dark blue), is quickly demoted to fifth and eventually drops out of the top-5 entirely due to Claude’s influence. Hepatitis B, initially ranked second by GPT-4 and top by Claude, stabilizes in the second position in rounds 3 and 4 (in light blue). Notably, Hepatitis C rises from second place on both lists to the top position and remains there (in black).
Figure 7: Convergence of all metrics
As demonstrated in the previous case study, Wasserstein distance (WD) effectively measures the divergence between LLM predictions and assesses debate convergence. Figures 6(b) and 6(c) show that WD stabilizes after three debate rounds, coinciding with a plateau in normalized mutual information (MI) between GPT-4 and Claude. This stabilization suggests their predictions converge.
Figure 7 illustrates the convergence of all divergence metrics—including Jensen-Shannon divergence, cross-entropy, and Kullback-Leibler divergence—particularly between the second and third rounds. Although the final joint prediction for Hepatitis C reached a high consensus of 37.5%, it deviates from the actual condition of Jaundice, which the Kaggle dataset reports with 10% confidence. EVINCE provides general practitioners with alerts and suggests remedial actions (see Appendices D.9 and C.8) to address this discrepancy. Recommended actions include querying additional symptoms from the patient and conducting specific laboratory tests.
4.4 Experiment Remarks
  • Suggestions to query symptom onset, duration, severity, trends, and associated symptoms (documented in Appendices B.8 and C.9).
5 Concluding Remarks
This paper introduces EVINCE, a framework designed to facilitate collaborative dynamics among Large Language Models (LLMs) through structured, adversarial debates. Our research demonstrates that EVINCE significantly advances the pursuit of Artificial General Intelligence (AGI) by enhancing three core characteristics: versatility, adaptivity, and reasoning capability. By addressing key limitations of current AI systems, including hallucination, bias, and limited reasoning, EVINCE provides a novel pathway towards more robust and capable artificial intelligence.

### v4

@@ 2.3 Maxims with Theoretical Foundations
Progress towards the optimality goal is guided and measured by metrics introduced in Section 2.1. This section explains how these metrics can be used in complementary ways to facilitate proper trade-offs between diversity and convergence, exploration and exploitation, and several other factors. In the $\mathsf{EVINCE}$ algorithm presented in Figure 1, we have annotated the steps to which these maxims are applied.
Maxim #1. Orchestrate Two Equally Competent LLMs in Structured Debate:
@@
Methods: Choosing LLMs with comparable performance on a shared validation set, a balanced debate can be ensured. Suitable models (in 2024) include GPT-4, Claude, and Gemini. Conditioning different instances of the same LLM to support opposing stances on a subject matter can also be effective due to the theoretical justification of in-context learning with conditional Bayesian statistics (Xie et al., 2021).
Maxim #2. Fostering Exploration through Diverse Perspectives:
While LLMs default to maximum likelihood predictions favoring popular outcomes, context conditioning can shift them toward exploration over exploitation. High initial contentiousness encourages dynamic debate and challenges to prevailing views, mitigating confirmation bias through contrary queries and diverse top-k predictions.
Methods: Set high initial contentiousness to stimulate exploration, enabling LLMs to challenge each other’s predictions. Monitor mutual information and agreement diversity to ensure productive debate.
Maxim #3. Refining High-Quality Perspectives:
Once new insights plateau (measured by mutual information, Wasserstein distance, and KL divergence), shift from exploration to exploitation by reducing contentiousness.
Methods: Algorithm Update($\Delta$) guides this transition to a conciliatory stage, focusing on strengthening well-supported arguments for the final output.
@@
Combine the probability distributions from two LLMs by weighting them based on distributional diversity and argument quality. While LLMs face fundamental challenges in generating accurate probabilities ((Peng & Zhang, 2024; Scholten et al., 2024)), we can use reasoning-calibrated confidence scores as an effective proxy for weighting predictions.
Methods: Following these four maxims:
  • Maxim #4.1 Prediction Reliability: Use entropy-based measures to estimate reliability. Lower entropy suggests higher confidence and greater reliability in predictions.
  • Maxim #4.2 Argument Quality: Evaluate argument quality using $\mathsf{CRIT}$, identifying logical fallacies and assessing the relevance and credibility of evidence.
  • Maxim #4.3 Aggregation: Apply a weighted aggregation method, such as a Bayesian model, to combine predictions, accounting for both probabilistic insights and argument quality.
  • Maxim #4.4 Diagnosis and RAG: If the final prediction has high entropy, indicating uncertainty, diagnose the issue and perform Retrieval-Augmented Generation (RAG) to obtain missing information.
@@ 2.4 Entropy Duality Theorem (EDT)
Please see Appendix C. ∎
3 Experiments on Healthcare & News
This study evaluates $\mathsf{EVINCE}$’ versatility by examining two distinct applications: 1. Disease Diagnosis, using LLMs as diagnostic tools, and 2. News Debiasing, which addresses bias in language generation. Together, these experiments demonstrate the framework’s capability to handle diverse and complex tasks.
@@ 3 Experiments on Healthcare & News
    Contentiousness & Convergence: The initial disagreement between the LLM (measured by the Wasserstein distance) reflects the diversity of perspectives, with disagreements decreasing as the debates progress and the LLMs reach consensus. Individual uncertainty (measured via Shannon entropy) follows a similar trend, starting high and reducing as LLMs refine their predictions.
@@
    EDT Effectiveness: LLM pairs optimized under the Entropy Duality Theorem (EDT) exhibit complementary error patterns. This synergy leads to improved overall prediction accuracy compared to non-EDT configurations, demonstrating $\mathsf{EVINCE}$’s ability to balance exploration and consensus.
@@
    Anomaly Identification: $\mathsf{EVINCE}$ enhances the detection of anomalies or inconsistencies in the generated predictions, uncovering ambiguities in input data or errors in existing labels in various applications.
General Problem Statement:
Given an input context $\kappa$ and a set of characteristics $F$, the objective is to generate and rank the top-$k$ predictions (e.g., diseases for the first problem and biased news for the second) from a set of possible outcomes $C$. This can be expressed as $P=\text{LLM}(F,\kappa)$, where each LLM outputs confidence scores for its top-$k$ predictions ($k\leq|C|$), leveraging the input features and context.
$P=\left(p(\text{top},1,\text{to},k\in C\mid F,\kappa)\right).\vspace{-0.1in}$ (1)
@@
Figure 2: EVINCE improves diagnosis accuracy markedly.
Context $\kappa$ enables dual entropy adjustment via temperature, top-$k$, and contentiousness $\Delta$, with higher settings increasing entropy and lower settings reducing it.
Note: Maxim #4 states that the initial confidence distribution $P$ from LLMs (Eq.1) represents confidence scores rather than absolute probabilities. In the $\mathsf{EVINCE}$ framework, these scores are calibrated through multi-iteration debates where predictions must be supported by arguments and examined by counterarguments. The $\mathsf{CRIT}$ algorithm (detailed in Appendix B) evaluates reasoning quality to adjust these confidence scores—reducing credibility when high-confidence predictions lack strong supporting evidence.
Since $\mathsf{EVINCE}$ and $\mathsf{CRIT}$ can employ different LLMs, and $\mathsf{CRIT}$ focuses solely on evaluating reasoning quality via Socratic methods, this collaborative approach uses reasoning quality as a proxy for confidence calibration. This sidesteps the fundamental challenge of generating reliable absolute probabilities from LLMs ((Peng & Zhang, 2024; Scholten et al., 2024)).
For example, if an LLM outputs confidence scores $P$ = (0.5, 0.3, 0.2) for its top-3 predicted diseases, and $\mathsf{CRIT}$ evaluates their supporting arguments as having strength (50%, 100%, 50%), the calibrated scores become (0.25, 0.3, 0.1). After normalization, the final distribution is $P$ = (0.38, 0.46, 0.16).
@@
3.1 Experiment #1 Disease Diagnosis
Resources & Dataset:
We use a Kaggle dataset (Patil, 2020) of 4,921 patient records, each containing a diagnosis and up to 17 symptoms. After deduplication, 304 unique cases across 40 diseases remain (dataset provided as supplement). We evaluate GPT-4, Gemini, and Claude-3 using their inherent knowledge, without few-shot training. Computing resources provided by Azure research grant.
Metric:
We use top-k Mean Reciprocal Rank (MRR): if a top-k prediction matches ground truth, the score is the reciprocal of its rank (1 for first, 1/2 for second, etc.); 0 if no match.
3.1.1 EVINCE vs. Baseline Accuracy
We evaluated $\mathsf{EVINCE}$ through pairwise collaborations of GPT-4, Gemini, and Claude-3 on 304 patient cases.
Parameter Settings:
Each LLM pair uses $k=5$ predictions with contrasting temperatures (high/low) and high contentiousness ($\Delta=0.9$). This ensures meaningful interaction while maximizing cross entropy for information exchange.
Evaluations:
We conducted two types of experiments:
1. Constrained Prediction (Baseline): Limiting predictions to the dataset’s 40 disease labels yielded high accuracy (95-97%). While unrealistic for general practice where physicians consider all possibilities, this constraint demonstrates LLMs’ flexibility in avoiding overfitting to potentially erroneous labels.
2. Unconstrained Prediction: Removing label constraints to simulate real-world conditions, all 304 cases showed consistent results across LLMs (standard deviation = 1.5%). Individual pre-debate accuracies (Figure 2, light blue bars) were: GPT-4 (82.8%), Gemini (80.3%), and Claude (79.5%).
@@ 3.1 Experiment #1 Disease Diagnosis
Performance:
$\mathsf{EVINCE}$ pairings (GPT-4/Claude-3 and GPT-4/Gemini-3) improved accuracy by 4-5 percentage points (Figure 2, green bars). The GPT-4/Claude-3 pairing achieved 87.5% accuracy (Figure 2(a)), comparable to state-of-the-art clinical algorithms (Peng et al., 2018).
Discussion:
While $\mathsf{EVINCE}$ significantly outperforms single LLM predictions, the remaining 12.5% inaccuracy merits deeper analysis. Given the reported 11% US misdiagnosis rate (Newman-Toker et al., 2023b), some discrepancies might reflect errors in the original dataset rather than $\mathsf{EVINCE}$ limitations. This observation suggests $\mathsf{EVINCE}$’s potential for identifying and correcting dataset errors, which we explore in Section 3.1.3.
3.1.2 Analysis: Explorative vs. Exploitative
(a) Study #2 convergence
(b) Study #3 convergence
Figure 5: Convergence of all information metrics.
@@ Discussion:
The confusion matrices for Hepatitis types A-E diagnosis (Fig. 3) reveal:
  • GPT-4: Limited accuracy (50%) for types A/B, poor performance on C/D.
  • Claude: Broader distribution of predictions of all types.
Claude’s higher entropy predictions complement GPT-4’s more confident but potentially incorrect assessments. This differential entropy drives the productive debate in $\mathsf{EVINCE}$, balancing exploration with exploitation to improve diagnostic accuracy.
@@
EVINCE achieves 5% higher precision in the diagnosis of specific types of liver disease compared to a baseline approach (Figure 2(a)), highlighting its ability to handle complex diagnostic scenarios.
Two primary factors contribute to the improved diagnostic accuracy of $\mathsf{EVINCE}$. First, structured debates with reasoning encourage LLMs to explore alternative diagnoses both broadly and deeply, leading to more comprehensive analysis and decision-making (refer to Appendices D and E for two debate examples). Second, pairing LLMs with high- and low-entropy prediction distributions, or those with a large Wasserstein distance (WD) between them, balances exploratory diversity with exploitative stability. This approach results in more robust and higher-quality decisions, as demonstrated in this second study.
@@
Observations from Information Metrics:
  • Entropy Stabilization: Figure 4(a) shows the stabilization of the entropy levels of both LLMs after three debate rounds, indicating convergence towards a similar stable entropy state.
  • Wasserstein Distance Improvement: Figure 4(b) shows a consistent improvement in the Wasserstein distance (WD) between the predictions of the two models over rounds.
  • Mutual Information Increase: Figure 4(c) reveals a 14% improvement in normalized mutual information between GPT-4 and Claude’s prediction distributions, suggesting increased shared information throughout the debate.
  • Convergence of Divergence Metrics: Figure 5(a) shows consistent convergence of all divergence metrics.
3.1.3 Explainability and Remediation
This study illustrates how $\mathsf{EVINCE}$ can identify potential misdiagnoses, explain the reasoning behind them, and recommend corrective actions. Traditionally, machine learning scientists rely on labeled data as “ground truth.” However, as evidenced by research like that of (Newman-Toker et al., 2023a) from Johns Hopkins, misdiagnosis is a widespread issue in healthcare systems globally. These erroneous diagnoses, often treated as ground truth, can be perpetuated by supervised learning algorithms, exacerbating the problem within the healthcare system. $\mathsf{EVINCE}$’s dialogue capabilities provide insights into the decision-making process and highlight missing information, helping to rectify erroneous predictions and redefine the ground truth. This approach offers a potential solution to the compounding effect of misdiagnoses in machine learning applications in healthcare.
Figure 6, which plots the respective entropies of GPT-4 and Claude, reveals two key insights. First, an large gap in Wasserstein Distance (WD) exists between the two models in the initial rounds. This disparity underscores the role of dual entropy in fostering information exchange. As the entropy values converge in rounds 3 and 4 and WD decreases significantly, we observe a corresponding convergence and stabilization of their mutual information. The information metrics that EVINCE employs effectively show the progress of the dialogue and convergence from explorative to consensus. Figure 5(b) illustrates the convergence of all divergence metrics—including Jensen-Shannon divergence, cross-entropy, and Kullback-Leibler divergence—particularly between the second and third rounds.
Appendix F demonstrates that EVINCE recommends additional symptoms to inquire about with the patient, as well as lab tests to confirm the diagnosis. These suggestions, validated by our hospital partners, provide valuable information to enhance diagnostic accuracy and correct errors.
3.2 Experiment #2 News Debiasing
This experimental framework aims to assess the feasibility of both detecting biases in textual content and implementing effective mitigation strategies. The first experiment focuses on bias detection, while the second explores the generation of balanced textual outputs as a corrective measure, moving beyond the limitations of prior studies that primarily focused only on identification.
Resources & Dataset:
This study utilizes a unique dataset of 619 news articles (54.3% about Democrat scandals, 45.7% about Republican scandals) selected from a larger 2013 repository of 14,033 articles compiled by fifteen reputable news organizations (Budak et al., 2016). These articles span diverse topics including civil rights, healthcare, elections, and national security, offering a comprehensive view of political coverage. Please visit (Anonymous, 2024) for links to the full set of news articles.
The dataset’s distinctive feature is its ground-truth labels provided by annotators with declared political affiliations. Through Amazon Mechanical Turk, 749 qualified U.S. workers, each annotating up to 1,000 randomly selected articles, classified articles on a five-point scale from ‘negatively biased’ to ‘positively biased’ (Budak et al., 2016). Crucially, each scandal article in our subset received independent classifications from both Democrat and Republican annotators. In Appendix G.1, we justify the sufficiency of the annotation method.
Metric:
The bias gap, which varies from -2 to 2, serves as a metric to quantify the direction and degree of bias. As a baseline, we used Claude and GPT-4 to generate initial results. For $\mathsf{EVINCE}$ experiments, we employed two instances of GPT-4, as Claude demonstrated a tendency to shift its predictions easily (a point elaborated on later).
3.2.1 Bias Evaluation Results
Due to space constraints, detailed results and discussions on Democrat scandals are provided in Appendix G.2, while those on Republican scandals are in Appendix G.3. Below, we summarize the key findings.
Figure 7 illustrates the distribution of ratings for all scandals across four scenarios:
1) Democrat-leaning annotators rating Democrat scandals, 2) Republican-leaning annotators rating Democrat scandals, 3) Democrat-leaning annotators rating Republican scandals, and 4) Republican-leaning annotators rating Republican scandals.
Figure 7: Bias Rating Distributions Show Strong Biases. D is more negative on how D scandals were reported (the sub-figure on the left), R is more negative on how R scandals were reported (the sub-figure on the right).
The figure reveals a clear pattern: Democrat-leaning annotators tend to rate news about Democrat scandals more negatively, while Republican-leaning annotators exhibit similar negativity towards reports on Republican scandals. The gap between these ratings is approximately one class-label (e.g., between “weak negative” and “neutral”), highlighting a tendency within both parties to defend their own and criticize the opposition.
$\mathsf{EVINCE}$, operating without emotional influence and refined through structured debate, consistently provides a more balanced, centrist perspective. This contributes to a more impartial discourse by mitigating partisan biases. $\mathsf{EVINCE}$’s justifications, documented in Appendix G, are transparent and reasonable. An editorial board can review these findings and decide whether to adjust labels or present both perspectives with explanations.
This experiment demonstrates that $\mathsf{EVINCE}$ effectively delivers centrist judgments supported by rationales. For a deeper understanding of $\mathsf{EVINCE}$’s bias assessment process, comprehensive justifications for each of the 31 analyzed articles are available in Appendix I.
3.2.2 Bias Mitigation Results
This experiment illustrates $\mathsf{EVINCE}$’s ability to identify bias in text, provide reasoned justifications, and propose remediation through the integration of diverse perspectives. We demonstrate how $\mathsf{EVINCE}$ utilizes statistical and information theory metrics to facilitate multi-agent dialogue, circumventing the “maximum likelihood” trap inherent in next-token generation and uncovering information from multiple viewpoints.
Figure 8: Convergence of all metrics, Wasserstein, normalized mutual information, normalized cross entropy
Using the example of the Euro-centric perspective on Christopher Columbus’ Wikipedia page regarding his voyages to America, $\mathsf{EVINCE}$ employs two GPT-4 instances: Agent A, supporting the Euro-centric view, and Agent B, opposing it. Table 5 in Appendix G.4 summarizes Agent A’s key arguments and its evolving stance by taking Agent B’s perspective into account, throughout the debate.
Guided by the maxims and entropy duality theorem from Section 2, we initiate the debate by prompting both agents to defend their positions rigorously and score each other’s bias using a five-label distribution (negative, weak negative, neutral, weak positive, positive). Figure 8 tracks the dialogue’s progress through Wasserstein distance (WD) (Kantorovich, 1942), normalized cross entropy (CE) (Shannon, 1948), and normalized mutual information (MI) (Cover & Thomas, 2006a). Initially, each agent is expected to perceive itself as neutral and the other as biased. The debate concludes when the bias distributions converge and mutual information plateaus, indicating a shared understanding.
4 Conclusion: Advancing Multi-Agent Debate
$\mathsf{EVINCE}$ not only enhances accuracy but also improves decision transparency and uncovers missing information, offering critical insights for error correction. Unlike prior debate-based methods that rely on redundancy, $\mathsf{EVINCE}$ enables structured information discovery, bias mitigation, and decision-making by balancing broad exploration with deep evaluative reasoning.